# The Equational Theories Project: Advancing Collaborative Mathematical Research at Scale

### v1

@@ 10.3. Using automated theorem provers
These steps eliminated all candidates of order less than 88, [^25] In particular we recover the nonexistence of laws of order $6$ characterizing Ward quasigroups, already announced by McCune and Kinyon at https://www.cs.unm.edu/~mccune/projects/gtsax/#division. and left only $213$ laws of order $8$ that could be equivalent to $\mathrm{E}_{\mathrm{HN}}$. These laws come in $31$ families consisting of a “parent” $5$-variable law and some specializations with pairs of variables being identified. The lowest-numbered law in this list is McCune’s law
@@ 11. AI and Machine Learning contributions
To model the implication graph, we used a convolutional neural network (CNN). For each pair of equations $(p,q)$, the input of the CNN consisted of a character-level tokenization (no vectorization) of the two equations, and the output of the CNN was a yes/no label depending on whether $p$ implies $q$. The CNN processed the input data using a 5-layer architecture, each layer composed of a 1-D convolution, followed by batch normalization and rectified linear unit activation functions [Goodfellow-et-al-2016]. After the last convolutional layer, a flattening layer and a softmax activation function were used to obtain the output of the network, i.e., the prediction of the implication for the input pair of equations. Note that we trained our models with the 2024-10-16 version of the (infinite) implication graph, for which 362 of the hardest implications (less than 0.002% of the total) were still unknown.
Prior to training the CNN, we divided the data into training (60%), validation (20%), and test (20%) subsets. The CNN was implemented in TensorFlow 2.9.0 [tensorflow2015-whitepaper] and trained on an NVIDIA 3080 Ti GPU with the following configuration: the binary cross-entropy as the loss function to minimize, the Adam method with an initial learning rate of $10^{-3}$ for the adjustment of network weights, a batch size of 1024 with random shuffling, a learning rate reduction by a factor of 2 after 15 epochs without improvement in the validation loss, and early stopping if no improvement occurred for 40 epochs, with the model with the lowest validation loss being retained as the final CNN. The final CNN was evaluated on the test set, reaching a prediction accuracy of 99.7%, which means that the model misclassified around 66k of the 22 million implications. Since this accuracy was somewhat surprising for such a small and “simple” model, as a control we generated a random label (yes/no) for each pair of equations, then trained the CNN on this data with 60%/20%/20% training/validation/test percentages, resulting in a 49.99% accuracy, as expected from an unbiased model.
@@
Labelled pairs of equations 1.5GB 9MB
Bit table 42MB 40KB
CNN (99.7% accuracy) 1.34MB 700KB
@@ 14. Conclusions and future directions
  • Low levels of required mathematical and formal prerequisites. The problems considered in the project did not require advanced mathematical knowledge (beyond a general familiarity with abstract algebra), nor a sophisticated understanding of formal proof assistants. This permitted contributions from a broad spectrum of participants, including those without a graduate mathematical training, as well as mathematicians with no experience in proof formalization. At a technical level, it also meant that formalization of proofs into Lean could be done immediately once certain base definitions (such as Magma) were constructed. This can be compared for instance with the recent formalization of the Polynomial Freiman–Ruzsa conjecture [^37] https://github.com/teorth/pfr, in which significant effort was expended in the first few days to settle on a suitable framework to formalize the mathematics of Shannon entropy. While some more sophisticated formal structures (such as the syntactic description of laws as pairs of words in a FreeMagma) were later introduced in the project, it was relatively straightforward to refactor previously written code to be compatible with these structures as they were incorporated into the project.
@@
    Does the law $\mathrm{x}\simeq\mathrm{y}\diamond(\mathrm{y}\diamond(\mathrm{y}\diamond(\mathrm{x}\diamond(\mathrm{z}\diamond\mathrm{y}))))$ (E5093) have any infinite models? In [Kisielewicz2] it was shown that it has no non-trivial finite models, but the infinite model case was left as an open question. A partial classification of laws of order $5$ with infinite models but no finite models is given at https://teorth.github.io/equational_theories/blueprint/order-5-austin-laws.html.
@@ Acknowledgments
We are grateful to the many additional participants in the Equational Theories Project for their numerous comments and encouragement, with particular thanks to Stanley Burris, Edward van de Meent and David Roberts. We warmly thank Michael Kinyon for generously sharing his expertise with Prover9–Mace4, and we are likewise grateful to Laura Kovács, Márton Hajdu, Martin Suda, and Michael Rawson of the Vampire development team for their helpful explanations regarding Vampire’s options and functionality. Additionally, we note that Shreyas Srinivas is a doctoral student at the Saarbrücken Graduate School for Computer Science.
@@ Appendix A Numbering system
We say that two equational laws are definitionally equivalent [^43] This can be distinguished from the weaker notion of propositional equivalence (mutual entailment) used in the rest of the paper. if one can be obtained from another by some combination of relabeling the variables and applying the symmetric law $w_{1}\simeq w_{2}\iff w_{2}\simeq w_{1}$. For instance, $(0\diamond 1)\diamond 2\simeq 1$ is definitionally equivalent to $0\simeq(1\diamond 0)\diamond 2$. We then replace every equational law with their minimal element in their definitional equivalence class, which can be viewed as the normal form for that law; for instance, the normal form of $(0\diamond 1)\diamond 2\simeq 1$ would be $0\simeq(1\diamond 0)\diamond 2$. Finally, we eliminate any law of the form $w\simeq w$ other than $0\simeq 0$. We then number the remaining equations $\mathrm{E}1,\mathrm{E}2,\dots$. For instance, $\mathrm{E}1$ is the trivial law $0\simeq 0$, $\mathrm{E}2$ is the constant law $0\simeq 1$, $\mathrm{E}3$ is the idempotent law $0\simeq 0\diamond 0$, and so forth. Lists and code for generating these equations, or the equation number attached to a given equation, can be found in the ETP repository.
@@
(E41) $\mathrm{x}\diamond\mathrm{x}$ $\simeq\mathrm{y}\diamond\mathrm{z}$ (Constant law)
@@ Appendix B Author contributions
  • Jose Brox: IMUVA-Mathematics Research Institute, Universidad de Valladolid, josebrox@uva.es. Supported by a postdoctoral fellowship “Convocatoria 2021” funded by Universidad de Valladolid, and partially supported by grant PID2022-137283NB-C22 funded by MCIN/AEI/10.13039/501100011033 and ERDF “A way of making Europe”
@@
  • Mario Carneiro: Chalmers University of Technology & Gothenburg University, Sweden, marioc@chalmers.se
  • Floris van Doorn: University of Bonn, vdoorn@math.uni-bonn.de
  • Martin Dvorak: Institute of Science and Technology Austria, martin.dvorak@matfyz.cz
  • Andrés Goens: TU Darmstadt, andres.goens@tu-darmstadt.de
  • Aaron Hill: Unaffiliated, aa1ronham@gmail.com
  • Harald Husum: Intelecy, harald.husum@intelecy.com
@@
  • Lorenzo Luccioli: University of Bologna, lorenzo.luccioli2@unibo.it
  • Douglas McNeil: Unaffiliated, dsm054@gmail.com
  • Alex Meiburg: Perimeter Institute for Theoretical Physics / University of Waterloo Institute for Quantum Computing, teqtp@ohaithe.re
  • Pietro Monticone: University of Trento, pietro.monticone@studenti.unitn.it
@@
  • Giovanni Paolini: University of Bologna, g.paolini@unibo.it
@@
  • Bernhard Reinke: Aix-Marseille Université, bernhard.reinke@univ-amu.fr
  • David Renshaw: Institute for Computer-Aided Reasoning in Mathematics, renshaw@icarm.io
  • Marcus Rossel: Barkhausen Institut, marcus.rossel@barkhauseninstitut.org
  • Cody Roux: Amazon Web Services, cody.roux@gmail.com
  • Jérémy Scanvic: Laboratoire de Physique, École Normale Supérieure de Lyon, jeremy.scanvic@ens-lyon.fr
  • Shreyas Srinivas: CISPA Helmholtz Center for Information Security, Saarbrücken, Germany, shreyas.srinivas@cispa.de
  • Anand Rao Tadipatri: University of Cambridge, art71@cam.ac.uk
@@
  • Fernando Vaquerizo-Villar: Biomedical Engineering Group, University of Valladolid, and CIBER de Bioingeniería, Biomateriales y Nanomedicina, Instituto de Salud Carlos III, fernando.vaquerizo@uva.es
  • Daniel Weber: Ben-Gurion University of the Negev, weberdan@post.bgu.ac.il

### v2

@@ 10.3. Using automated theorem provers
These steps eliminated all candidates of order less than 88, [^23] In particular we recover the nonexistence of laws of order $6$ characterizing Ward quasigroups, already announced by McCune and Kinyon [46]. and left only $213$ laws of order $8$ that could be equivalent to $\mathrm{E}_{\mathrm{HN}}$. These laws come in $31$ families consisting of a “parent” $5$-variable law and some specializations with pairs of variables being identified. The lowest-numbered law in this list is McCune’s law
@@ 11. AI and Machine Learning contributions
To model the implication graph, we used a convolutional neural network (CNN). For each pair of equations $(p,q)$, the input of the CNN consisted of a character-level tokenization (no vectorization) of the two equations, and the output of the CNN was a yes/no label depending on whether $p$ implies $q$. The CNN processed the input data using a 5-layer architecture, each layer composed of a 1-D convolution, followed by batch normalization and rectified linear unit activation functions [27]. After the last convolutional layer, a flattening layer and a softmax activation function were used to obtain the output of the network, i.e., the prediction of the implication for the input pair of equations. Note that we trained our models with the 16 October 2024 version of the (infinite) implication graph, for which 362 of the hardest implications (less than 0.002% of the total) were still unknown.
Prior to training the CNN, we divided the data into training (60%), validation (20%), and test (20%) subsets. The CNN was implemented in TensorFlow 2.9.0 [1] and trained on an NVIDIA 3080 Ti GPU with the following configuration: the binary cross-entropy as the loss function to minimize, the Adam method with an initial learning rate of $10^{-3}$ for the adjustment of network weights, a batch size of 1024 with random shuffling, a learning rate reduction by a factor of 2 after 15 epochs without improvement in the validation loss, and early stopping if no improvement occurred for 40 epochs, with the model with the lowest validation loss being retained as the final CNN. The final CNN was evaluated on the test set, reaching a prediction accuracy of 99.7%, which means that the model misclassified around 66k of the 22 million implications. Since this accuracy was somewhat surprising for such a small and “simple” model, as a control we generated a random label (yes/no) for each pair of equations, then trained the CNN on this data with 60%/20%/20% training/validation/test percentages, resulting in a 49.99% accuracy, as expected from an unbiased model.
@@
Labelled pairs of equations $1.5\text{\,}\mathrm{GB}$ $9\text{\,}\mathrm{MB}$
Bit table $42\text{\,}\mathrm{MB}$ $40\text{\,}\mathrm{kB}$
CNN (99.7% accuracy) $1.34\text{\,}\mathrm{MB}$ $700\text{\,}\mathrm{kB}$
@@ 14. Conclusions and future directions
  • Low levels of required mathematical and formal prerequisites. The problems considered in the project did not require advanced mathematical knowledge (beyond a general familiarity with abstract algebra), nor a sophisticated understanding of formal proof assistants. This permitted contributions from a broad spectrum of participants, including those without a graduate mathematical training, as well as mathematicians with no experience in proof formalization. At a technical level, it also meant that formalization of proofs into Lean could be done immediately once certain base definitions (such as Magma) were constructed. This can be compared for instance with the recent formalization of the Polynomial Freiman–Ruzsa conjecture [22], in which significant effort was expended in the first few days to settle on a suitable framework to formalize the mathematics of Shannon entropy. While some more sophisticated formal structures (such as the syntactic description of laws as pairs of words in a FreeMagma) were later introduced in the project, it was relatively straightforward to refactor previously written code to be compatible with these structures as they were incorporated into the project.
@@
    Does the law $\mathrm{x}\simeq\mathrm{y}\diamond(\mathrm{y}\diamond(\mathrm{y}\diamond(\mathrm{x}\diamond(\mathrm{z}\diamond\mathrm{y}))))$ (E5093) have any infinite models? In [34] it was shown that it has no non-trivial finite models, but the infinite model case was left as an open question. A partial classification of laws of order $5$ with infinite models but no finite models is given at https://teorth.github.io/equational_theories/blueprint/order-5-austin-laws.html.
@@ Acknowledgments
We are grateful to the many additional participants in the Equational Theories Project for their numerous comments and encouragement, with particular thanks to Stanley Burris, Edward van de Meent and David Michael Roberts. We warmly thank Michael Kinyon for generously sharing his expertise with Prover9-Mace4, and we are likewise grateful to Laura Kovács, Márton Hajdu, Martin Suda, and Michael Rawson of the Vampire development team for their helpful explanations regarding Vampire’s options and functionality. We gratefully acknowledge the computational resources provided by the SINGACOM research group at IMUVa, the Mathematics Research Institute of the University of Valladolid. Additionally, we note that Shreyas Srinivas is a doctoral student at the Saarbrücken Graduate School for Computer Science.
@@ Appendix A Numbering system
We say that two equational laws are definitionally equivalent [^40] This can be distinguished from the weaker notion of propositional equivalence (mutual entailment) used in the rest of the paper. if one can be obtained from another by some combination of relabeling the variables and applying the symmetric law $w_{1}\simeq w_{2}\iff w_{2}\simeq w_{1}$. For instance, $(0\diamond 1)\diamond 2\simeq 1$ is definitionally equivalent to $0\simeq(1\diamond 0)\diamond 2$. We then replace every equational law with their minimal element in their definitional equivalence class, which can be viewed as the normal form for that law; for instance, the normal form of $(0\diamond 1)\diamond 2\simeq 1$ would be $0\simeq(1\diamond 0)\diamond 2$. Finally, we eliminate any law of the form $w\simeq w$ other than $0\simeq 0$. We then number the remaining equations $\mathrm{E}1,\mathrm{E}2,\dots$. For instance, $\mathrm{E}1$ is the trivial law $0\simeq 0$, $\mathrm{E}2$ is the singleton law $0\simeq 1$, $\mathrm{E}3$ is the idempotent law $0\simeq 0\diamond 0$, and so forth. Lists and code for generating these equations, or the equation number attached to a given equation, can be found in the ETP repository.
@@
(E14) $\mathrm{x}$ $\simeq\mathrm{y}\diamond(\mathrm{x}\diamond\mathrm{y})$ (Semi-symmetric quasigroups)
@@
(E41) $\mathrm{x}\diamond\mathrm{x}$ $\simeq\mathrm{y}\diamond\mathrm{z}$ (Equivalent to (E46))
@@
(E450) $\mathrm{x}$ $\simeq\mathrm{x}\diamond(\mathrm{y}\diamond(\mathrm{z}\diamond(\mathrm{y}\diamond\mathrm{x})))$
(E492) $\mathrm{x}$ $\simeq\mathrm{y}\diamond(\mathrm{x}\diamond(\mathrm{z}\diamond(\mathrm{z}\diamond\mathrm{y})))$ (Totally symmetric quasigroups)
@@
(E650) $\mathrm{x}$ $\simeq\mathrm{x}\diamond(\mathrm{y}\diamond((\mathrm{z}\diamond\mathrm{x})\diamond\mathrm{y}))$ (Non-trivially equivalent to (E4))
@@
(E1323) $\mathrm{x}$ $\simeq\mathrm{y}\diamond(((\mathrm{y}\diamond\mathrm{y})\diamond\mathrm{x})\diamond\mathrm{y})$
@@
(E1518) $\mathrm{x}$ $\simeq(\mathrm{y}\diamond\mathrm{y})\diamond(\mathrm{x}\diamond(\mathrm{y}\diamond\mathrm{x}))$
@@
(E2744) $\mathrm{x}$ $\simeq((\mathrm{y}\diamond\mathrm{y})\diamond(\mathrm{y}\diamond\mathrm{x}))\diamond\mathrm{y}$
@@ Appendix B Author contributions
  • Jose Brox: IMUVA-Mathematics Research Institute, Universidad de Valladolid, josebrox@uva.es, ORCID 0000-0001-9822-5838. Supported by a postdoctoral fellowship “Convocatoria 2021” funded by Universidad de Valladolid, and partially supported by grant PID2022-137283NB-C22 funded by MCIN/AEI/10.13039/501100011033 and ERDF “A way of making Europe”
@@
  • Mario Carneiro: Chalmers University of Technology & Gothenburg University, Sweden, marioc@chalmers.se, ORCID 0000-0002-0470-5249
  • Floris van Doorn: University of Bonn, vdoorn@math.uni-bonn.de, ORCID 0000-0003-2899-8565
  • Martin Dvorak: Institute of Science and Technology Austria, martin.dvorak@matfyz.cz, ORCID 0000-0001-5293-214X
  • Andrés Goens: TU Darmstadt, andres.goens@tu-darmstadt.de, ORCID 0000-0002-0409-1363
  • Aaron Hill: Unaffiliated, aa1ronham@gmail.com, ORCID 0009-0007-0827-1277
  • Harald Husum: Intelecy, harald.husum@intelecy.com, ORCID 0009-0003-0634-7435
@@
  • Lorenzo Luccioli: University of Bologna, lorenzo.luccioli2@unibo.it, ORCID 0009-0009-2256-2833
  • Douglas McNeil: Unaffiliated, dsm054@gmail.com, ORCID 0009-0006-4662-0469
  • Alex Meiburg: Perimeter Institute for Theoretical Physics / University of Waterloo Institute for Quantum Computing, teqtp@ohaithe.re, ORCID 0000-0002-4506-9146
  • Pietro Monticone: University of Trento, pietro.monticone@studenti.unitn.it, ORCID 0000-0002-2731-9623
@@
  • Giovanni Paolini: University of Bologna, g.paolini@unibo.it, ORCID 0000-0002-3964-9101
@@
  • Bernhard Reinke: Aix-Marseille Université, bernhard.reinke@univ-amu.fr, ORCID 0000-0001-9024-2449
  • David Renshaw: Institute for Computer-Aided Reasoning in Mathematics, renshaw@icarm.io, ORCID 0000-0002-9987-9144
  • Marcus Rossel: Barkhausen Institut, marcus.rossel@barkhauseninstitut.org, ORCID 0009-0001-3567-6890
  • Cody Roux: Amazon Web Services, cody.roux@gmail.com, ORCID 0009-0004-5304-037X
  • Jérémy Scanvic: Laboratoire de Physique, École Normale Supérieure de Lyon, jeremy.scanvic@ens-lyon.fr, ORCID 0009-0003-6117-0492
  • Shreyas Srinivas: CISPA Helmholtz Center for Information Security, Saarbrücken, Germany, shreyas.srinivas@cispa.de, ORCID 0000-0002-3993-1596
  • Anand Rao Tadipatri: University of Cambridge, art71@cam.ac.uk, ORCID 0009-0007-0057-4169
@@
  • Fernando Vaquerizo-Villar: Biomedical Engineering Group, University of Valladolid, and CIBER de Bioingeniería, Biomateriales y Nanomedicina, Instituto de Salud Carlos III, fernando.vaquerizo@uva.es, ORCID 0000-0002-5898-2006
  • Daniel Weber: Ben-Gurion University of the Negev, weberdan@post.bgu.ac.il, ORCID 0009-0008-4615-6445